OpenStudy (anonymous):

solve the equation By Factoring: (2x+3)(3x-1)=14x

OpenStudy (lalaly):

\[\huge{6x^2-2x+9x-3}=14x\]

OpenStudy (lalaly):

\[\large{6x^2+7x-3=14x}\]

OpenStudy (lalaly):

\[\large{6x^2-7x-3=0}\]

OpenStudy (lalaly):

l\[\huge{(2x-3)(3x+1)}=0\]

OpenStudy (lalaly):

\[\large{x=\frac{3}{2}}, \large{x=\frac{-1}{3}}\]